# Optimal Transport for Measures with Noisy Tree Metric

### v1

@@ Abstract
We study optimal transport (OT) problem for probability measures supported on a tree metric space. It is known that such OT problem (i.e., tree-Wasserstein (TW)) admits a closed-form expression, but depends fundamentally on the underlying tree structure over supports of input measures. In practice, the given tree structure may be, however, perturbed due to noisy or adversarial measurements. In order to mitigate this issue, we follow the max-min robust OT approach which considers the maximal possible distances between two input measures over an uncertainty set of tree metrics. In general, this approach is hard to compute, even for measures supported in $1$-dimensional space, due to its non-convexity and non-smoothness which hinders its practical applications, especially for large-scale settings. In this work, we propose novel uncertainty sets of tree metrics from the lens of edge deletion/addition which covers a diversity of tree structures in an elegant framework. Consequently, by building upon the proposed uncertainty sets, and leveraging the tree structure over supports, we show that the max-min robust OT also admits a closed-form expression for a fast computation as its counterpart standard OT (i.e., TW). Furthermore, we demonstrate that the max-min robust OT satisfies the metric property and is negative definite. We then exploit its negative definiteness to propose positive definite kernels and test them in several simulations on various real-world datasets on document classification and topological data analysis for measures with noisy tree metric.
@@ 1 Introduction
Optimal transport (OT) has become a popular approach for comparing probability measures. OT provides a set of powerful tools that can be utilized in various research fields such as machine learning (Peyré & Cuturi, 2019; Nadjahi et al., 2019; Titouan et al., 2019; Bunne et al., 2019, 2022; Janati et al., 2020; Muzellec et al., 2020; Paty et al., 2020; Mukherjee et al., 2021; Altschuler et al., 2021; Fatras et al., 2021; Scetbon et al., 2021; Si et al., 2021; Takezawa et al., 2022; Fan et al., 2022), statistics (Mena & Niles-Weed, 2019; Weed & Berthet, 2019; Liu et al., 2022; Nguyen et al., 2022; Nietert et al., 2022; Wang et al., 2022), or computer graphics (Rabin et al., 2011; Solomon et al., 2015; Lavenant et al., 2018).
Following the recent line of research on leveraging tree structure to scale up OT problems (Le et al., 2019; Sato et al., 2020; Le & Nguyen, 2021; Takezawa et al., 2022; Yamada et al., 2022), in this work, we study OT problem for probability measures supported on a tree metric space. Such OT problem (i.e., tree-Wasserstein (TW)) not only admits a closed-form expression, generalizes the sliced Wasserstein (SW) [^1] SW projects supports into a $1$-dimensional space and exploits the closed-form expression of the univariate OT. (Rabin et al., 2011) (i.e., a tree is a chain), but also alleviates the limited capacity issue of SW to capture topological structure of input measures, especially in high-dimensional spaces, since it provides more flexibility and degrees of freedom by choosing a tree rather than a line (Le et al., 2019). However, it depends fundamentally on the underlying tree structure over supports of input measures. Nevertheless, in practical applications, the given tree structure may be perturbed due to noisy or adversarial measurements. For examples, (i) edge lengths may be noisy; (ii) for a physical tree, node positions may be perturbed or under adversarial attacks; (iii) some connecting nodes may be merged into each other; or (iv) some nodes may be duplicated and their corresponding edge lengths are positive.
For OT problem with noisy ground cost, a common approach in the literature is to consider the maximal possible distance between two input measures over an uncertainty set of ground metrics, i.e., the max-min robust OT (Paty & Cuturi, 2019; Deshpande et al., 2019; Lin et al., 2020). However, such approach usually leads to optimization problems which are challenging to compute due to their non-convexity and non-smoothness (Paty & Cuturi, 2019; Lin et al., 2020), even for input measures supported in $1$-dimensional spaces (Deshpande et al., 2019)). Another approach instead considers the min-max robust OT which is a convexified relaxation and is an upper bound of the max-min robust OT (Alvarez-Melis et al., 2018; Genevay et al., 2018; Paty & Cuturi, 2019; Dhouib et al., 2020).
@@
We emphasize both max-min and min-max robust OT approaches have their own advantages for the OT problem with noisy ground metric. In this work, we focus on the max-min robust OT for measures with a noisy tree metric ground cost. [^2] For general applications, one can sample tree metric for measures with supports in Euclidean space (see (Le et al., 2019)). At a high level, our main contributions are three-fold as follows:
@@
The paper is organized as follows: we give a brief recap of OT with tree metric cost in §2. In §3, we propose novel uncertainty sets of tree metrics, and leverage them to derive a closed-form expression for the max-min robust OT for measures with noisy tree metric. We show that it satisfies metric property and is negative definite. Consequently, we propose positive definite kernels built upon the robust OT. In §4, we discuss related work. In §5, we evaluate the proposed robust OT kernels for measures with noisy tree metric on document classification and TDA, and conclude our work in §6. Detailed proofs for our theoretical results are placed in the supplementary (§B).
Notations. We write $\mathbbm{1}$ for the vector of ones, and use $|E|$ to denote the cardinality of set $E$. For $1\leq p\leq\infty$, its conjugate is denoted by $p^{\prime}$, i.e., $p^{\prime}\in[1,\infty]$ s.t. $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. In particular, $p^{\prime}=\infty$ when $p=1$, and $p^{\prime}=1$ when $p=\infty$. Let $\left\lVert\cdot\right\rVert_{p}$ represent the $\ell_{p}$-norm in $\mathbb{R}^{|E|}$, and ¯$\overline{\mathcal{B}}_{p}(v,\lambda)\triangleq\{u\in\mathbb{R}^{|E|}:\|u-v\|_{p}\leq\lambda\}$ be the closed $\ell_{p}$-ball centering at $v\in\mathbb{R}^{|E|}$ and with radius $\lambda>0$. We denote $\delta_{x}$ as the Dirac function at $x$.
2 A Recap of Optimal Transport with Tree Metric Cost
In this section, we give a brief recap of OT with tree metric cost. We refer the readers to (Le et al., 2019) and the supplementary (§A.2–§A.3) for further details.
Tree metric. Let ${\mathcal{T}}=(V,E)$ be a tree rooted at $r$ with nonnegative weights $\{w_{e}\}_{e\in E}$ (i.e., edge length), where $V$ and $E$ are the sets of vertices and edges respectively. For any two nodes $x,z\in V$, we write $[x,z]$ for the unique path on ${\mathcal{T}}$ connecting $x$ and $z$. For an edge $e$, $\gamma_{e}$ denotes the set of all nodes $x$ such that the path $[r,x]$ contains the edge $e$. That is,
@@ 2 A Recap of Optimal Transport with Tree Metric Cost
Optimal transport (OT). For probability measures $\mu,\,\nu\in{\mathcal{P}}(V)$, let ${\mathcal{R}}(\mu,\nu)$ be the set of measures $\pi$ on the product space $V\times V$ such that $\pi(A\times V)=\mu(A)$ and $\pi(V\times B)=\nu(B)$ for all Borel sets $A,B\subset V$. By using tree metric $d_{{\mathcal{T}}}$ as the ground cost, the $1$-Wasserstein distance ${\mathcal{W}}_{{\mathcal{T}}}$ between $\mu$ and $\nu$ is defined as follows:
@@ 3 Robust Optimal Transport for Measures with Noisy Tree Metric
Max-min robust OT. Let $\mathbb{U}({\mathcal{T}})$ denote a family of tree metrics for the given tree ${\mathcal{T}}$. By considering $\mathbb{U}({\mathcal{T}})$ as the uncertainty set of tree metrics, the max-min robust OT between two input probability measures $\mu,\nu\in{\mathcal{P}}(V)$ is defined as follows:
@@
Uncertainty sets of tree metrics. Our key observation is that the computation of OT between input measures with tree metric does not depend on edges which have 00-length (i.e., $w_{e}=0$). We formally summarize it in Theorem 3.1.
@@
Let consider tree ${\mathcal{T}}_{1}$ as in Figure 1 and set $w_{e_{1}}=0$. Tree ${\mathcal{T}}_{2}$ is constructed from ${\mathcal{T}}_{1}$ as in Figure 1. Given two probability measures $\mu=a_{1}\delta_{r}+a_{2}\delta_{v_{1}}+a_{3}\delta_{v_{5}}$ and $\nu=b_{1}\delta_{v_{2}}+b_{2}\delta_{v_{4}}$ in ${\mathcal{T}}_{1}$ (i.e., $a_{1},a_{2},a_{3},b_{1},b_{2}\geq 0;a_{1}+a_{2}+a_{3}=1;b_{1}+b_{2}=1$), their corresponding measures in ${\mathcal{T}}_{2}$ are $\mu_{{\mathcal{T}}_{2}}=(a_{1}+a_{2})\delta_{r}+a_{3}\delta_{v_{5}}$ and $\nu_{{\mathcal{T}}_{2}}=\nu$ respectively. Then, we have ${\mathcal{W}}_{{\mathcal{T}}_{1}}(\mu,\nu)={\mathcal{W}}_{{\mathcal{T}}_{2}}(\mu_{{\mathcal{T}}_{2}},\nu_{{\mathcal{T}}_{2}})$.
@@
Let consider tree ${\mathcal{T}}_{1}$ as in Figure 1. Tree ${\mathcal{T}}_{3}$ is constructed from ${\mathcal{T}}_{1}$ as in Figure 1, and we set $w_{e_{a}}=w_{e_{b}}=0$. Given two probability measures $\mu=a_{1}\delta_{v_{1}}+a_{2}\delta_{v_{5}}$ and $\nu=b_{1}\delta_{v_{2}}+b_{2}\delta_{v_{4}}$ in ${\mathcal{T}}_{1}$ (i.e., $a_{1},a_{2},b_{1},b_{2}\geq 0;a_{1}+a_{2}=1;b_{1}+b_{2}=1$), their corresponding measures in ${\mathcal{T}}_{3}$ are $\mu_{{\mathcal{T}}_{3}}=a_{11}\delta_{v_{a}}+a_{12}\delta_{v_{1}}+a_{2}\delta_{v_{5}}$ where $a_{11},a_{12}\geq 0;a_{11}+a_{12}=a_{1}$ and $\nu_{{\mathcal{T}}_{3}}=\nu$ respectively. Then, regardless any split mass ($a_{11},a_{12}$) in $\mu_{{\mathcal{T}}_{3}}$ in ${\mathcal{T}}_{3}$ from original $\mu$ in ${\mathcal{T}}$, we have ${\mathcal{W}}_{{\mathcal{T}}_{1}}(\mu,\nu)={\mathcal{W}}_{{\mathcal{T}}_{3}}(\mu_{{\mathcal{T}}_{3}},\nu_{{\mathcal{T}}_{3}})$.
Therefore, we can add or delete edges with 00-length on the given tree ${\mathcal{T}}$ without changing the OT distance. More concretely, (i) for edge deletion, we collapse edges with 00-length in ${\mathcal{T}}$ by merging the two corresponding vertices of those edges together (see Example 3.2); (ii) for edge addition, we duplicate any vertex in ${\mathcal{T}}$ and connect them with 00-length edge (see Example 3.3). These actions help to transform the given tree structure ${\mathcal{T}}$ into various tree structures, which play the fundamental role to construct uncertainty sets with diverse tree structures. Additionally, we further vary edge lengths of these tree structures to derive our novel uncertainty sets of tree metrics.
@@
(i) Constraints on individual edge. We consider an uncertainty for each edge length $\hat{w}_{e}$ of edge $e\in E$ in tree ${\mathcal{T}}(\hat{w})$. Specifically, we consider $\hat{w}_{e}$ belongs to some uncertainty interval $w_{e}-\alpha_{e}\leq\hat{w}_{e}\leq w_{e}+\beta_{e}$ around the edge weight $w_{e}$ in tree ${\mathcal{T}}$. In the vector form, this just means that $w-\alpha\leq\hat{w}\leq w+\beta$, with $\alpha,\,\beta\in\mathbb{R}^{|E|}_{+}$ satisfying $\alpha\leq w$ (i.e., edge weights are nonnegative). Thus, $w-\alpha$ and $w+\beta$ are respectively the lower and upper limits of the uncertainty interval for the vector of edge lengths.
@@
Closed-form expressions. By building upon the proposed uncertainty sets of tree metrics and leveraging tree structure, we derive closed-form expressions for the max-min robust OT, similar to its counterpart standard OT (i.e., TW).
@@
Connection between two approaches. We next draw a connection between two approaches for the robust OT for measures with noisy tree metric as follows:
@@
Assume that $\beta=\lambda\mathbbm{1}$. Then, we have
@@
Computational complexity. From the closed-form expressions for $\textsf{RT}_{{\mathcal{U}}}$ in (7) and for $\textsf{RT}_{{\mathcal{U}}_{p}}$ in (10), the computational complexity of robust OT $\textsf{RT}_{{\mathcal{U}}}$ and $\textsf{RT}_{{\mathcal{U}}_{p}}$ for measures with noisy tree metric is linear to the number of edges in ${\mathcal{T}}$ (i.e., $\mathcal{O}(|E|)$), which is in the same order of computational complexity as the counterpart standard OT for measures with tree metric (i.e., TW) (Ba et al., 2011; Le et al., 2019). Recall that, in general, the max-min robust OT problem is hard and expensive to compute due to its non-convexity and non-smoothness (Paty & Cuturi, 2019; Lin et al., 2020), even for measures supported in $1$-dimensional space (Deshpande et al., 2019). [^4] Even with a given optimal ground metric cost, the computational complexity of max-min/min-max robust OT is in the same order as their counterpart standard OT (i.e., their objective function).
@@
Negative definiteness. We next prove that the robust OT for measures with noisy tree metric is negative definite. Therefore, we can derive positive definite kernels built upon the robust OT.
@@
Metric property. We end this section by showing that the robust OT is a metric.

### v2

@@ Abstract
We study optimal transport (OT) problem for probability measures supported on a tree metric space. It is known that such OT problem (i.e., tree-Wasserstein (TW)) admits a closed-form expression, but depends fundamentally on the underlying tree structure over supports of input measures. In practice, the given tree structure may be, however, perturbed due to noisy or adversarial measurements. To mitigate this issue, we follow the max-min robust OT approach which considers the maximal possible distances between two input measures over an uncertainty set of tree metrics. In general, this approach is hard to compute, even for measures supported in one-dimensional space, due to its non-convexity and non-smoothness which hinders its practical applications, especially for large-scale settings. In this work, we propose novel uncertainty sets of tree metrics from the lens of edge deletion/addition which covers a diversity of tree structures in an elegant framework. Consequently, by building upon the proposed uncertainty sets, and leveraging the tree structure over supports, we show that the robust OT also admits a closed-form expression for a fast computation as its counterpart standard OT (i.e., TW). Furthermore, we demonstrate that the robust OT satisfies the metric property and is negative definite. We then exploit its negative definiteness to propose positive definite kernels and test them in several simulations on various real-world datasets on document classification and topological data analysis.
@@ 1 INTRODUCTION
Optimal transport (OT) has become a popular approach for comparing probability measures. OT provides a set of powerful tools that can be utilized in various research fields such as machine learning (Peyré & Cuturi, 2019; Nadjahi et al., 2019; Titouan et al., 2019; Bunne et al., 2019, 2022; Janati et al., 2020; Muzellec et al., 2020; Paty et al., 2020; Mukherjee et al., 2021; Altschuler et al., 2021; Fatras et al., 2021; Scetbon et al., 2021; Si et al., 2021; Le et al., 2021a, b; Liu et al., 2021; Nguyen et al., 2021c; Takezawa et al., 2022; Fan et al., 2022; Hua et al., 2023; Nguyen & Ho, 2023), statistics (Mena & Niles-Weed, 2019; Weed & Berthet, 2019; Liu et al., 2022; Nguyen et al., 2022; Nietert et al., 2022; Wang et al., 2022; Pham et al., 2024), or computer vision/graphics (Rabin et al., 2011; Solomon et al., 2015; Lavenant et al., 2018; Nguyen et al., 2021b; Saleh et al., 2022).
Following the recent line of research on leveraging tree structure to scale up OT problems (Le et al., 2019; Sato et al., 2020; Le & Nguyen, 2021; Takezawa et al., 2022; Yamada et al., 2022), in this work, we study OT problem for probability measures supported on a tree metric space. Such OT problem (i.e., tree-Wasserstein (TW)) not only admits a closed-form expression, generalizes the sliced Wasserstein (SW) [^1] SW projects supports into a one-dimensional space and exploits the closed-form expression of the univariate OT. (Rabin et al., 2011) (i.e., a tree is a chain), but also alleviates the limited capacity issue of SW to capture topological structure of input measures, especially in high-dimensional spaces, since it provides more flexibility and degrees of freedom by choosing a tree rather than a line (Le et al., 2019). However, it depends fundamentally on the underlying tree structure over supports of input measures. Nevertheless, in practical applications, the given tree structure may be perturbed due to noisy or adversarial measurements. For examples, (i) edge lengths may be noisy; (ii) for a physical tree, node positions may be perturbed or under adversarial attacks; (iii) some connecting nodes may be merged into each other; or (iv) some nodes may be duplicated and their corresponding edge lengths are positive.
For OT problem with noisy ground cost, a common approach in the literature is to consider the maximal possible distance between two input measures over an uncertainty set of ground metrics, i.e., the max-min robust OT (Paty & Cuturi, 2019; Deshpande et al., 2019; Lin et al., 2020). However, such approach usually leads to optimization problems which are challenging to compute due to their non-convexity and non-smoothness (Paty & Cuturi, 2019; Lin et al., 2020), even for input measures supported in one-dimensional spaces (Deshpande et al., 2019)). Another approach instead considers the min-max robust OT which is a convexified relaxation and is an upper bound of the max-min robust OT (Alvarez-Melis et al., 2018; Genevay et al., 2018; Paty & Cuturi, 2019; Dhouib et al., 2020).
@@
We emphasize both max-min and min-max robust OT approaches have their own advantages for the OT problem with noisy ground metric. In this work, we focus on the max-min robust OT for measures with a noisy tree metric ground cost. [^2] One can sample tree metric for measures with supports in Euclidean space (see (Le et al., 2019)) to leverage TW for general applications, especially large-scale setting. At a high level, our main contributions are three-fold as follows:
@@
The paper is organized as follows: we give a brief recap of OT with tree metric cost in §2. In §3, we propose novel uncertainty sets of tree metrics, and leverage them to derive a closed-form expression for the max-min robust OT for measures with noisy tree metric. We show that it satisfies metric property and is negative definite. Consequently, we propose positive definite kernels built upon the robust OT. In §4, we discuss related work. In §5, we evaluate the proposed robust OT kernels for measures with noisy tree metric on document classification and TDA, and conclude our work in §6. Detailed proofs for our theoretical results are placed in the supplementary (§B). Additionally, we have released code for our proposed approach. [^4] https://github.com/lttam/RobustOT-NoisyTreeMetric
Notations. We write $\mathds{1}$ for the vector of ones, and use $|E|$ to denote the cardinality of set $E$. For $1\leq p\leq\infty$, its conjugate is denoted by $p^{\prime}$, i.e., $p^{\prime}\in[1,\infty]$ s.t. $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. In particular, $p^{\prime}=\infty$ when $p=1$, and $p^{\prime}=1$ when $p=\infty$. Let $\left\lVert\cdot\right\rVert_{p}$ represent the $\ell_{p}$-norm in $\mathbb{R}^{|E|}$, and ¯$\overline{\mathcal{B}}_{p}(v,\lambda)\triangleq\{u\in\mathbb{R}^{|E|}:\|u-v\|_{p}\leq\lambda\}$ be the closed $\ell_{p}$-ball centering at $v\in\mathbb{R}^{|E|}$ and with radius $\lambda>0$. We denote $\delta_{x}$ as the Dirac function at $x$.
2 A REVIEW OF OPTIMAL TRANSPORT WITH TREE METRIC COST
In this section, we give a brief recap of OT with tree metric cost. We refer the readers to Le et al. (2019) and the supplementary (§A.2–§A.3) for further details.
Tree metric.
Let ${\mathcal{T}}=(V,E)$ be a tree rooted at $r$ with nonnegative weights $\{w_{e}\}_{e\in E}$ (i.e., edge length), where $V$ and $E$ are the sets of vertices and edges respectively. For any two nodes $x,z\in V$, we write $[x,z]$ for the unique path on ${\mathcal{T}}$ connecting $x$ and $z$. For an edge $e$, $\gamma_{e}$ denotes the set of all nodes $x$ such that the path $[r,x]$ contains the edge $e$. That is,
@@ 2 A REVIEW OF OPTIMAL TRANSPORT WITH TREE METRIC COST
Optimal transport (OT).
For probability measures $\mu,\,\nu\in{\mathcal{P}}(V)$, let ${\mathcal{R}}(\mu,\nu)$ be the set of measures $\pi$ on the product space $V\times V$ such that $\pi(A\times V)=\mu(A)$ and $\pi(V\times B)=\nu(B)$ for all Borel sets $A,B\subset V$. By using tree metric $d_{{\mathcal{T}}}$ as the ground cost, the $1$-Wasserstein distance ${\mathcal{W}}_{{\mathcal{T}}}$ between $\mu$ and $\nu$ is defined as follows:
@@ 3 ROBUST OPTIMAL TRANSPORT FOR MEASURES WITH NOISY TREE METRIC
Max-min robust OT.
Let $\mathbb{U}({\mathcal{T}})$ denote a family of tree metrics for the given tree ${\mathcal{T}}$. By considering $\mathbb{U}({\mathcal{T}})$ as the uncertainty set of tree metrics, the max-min robust OT between two input probability measures $\mu,\nu\in{\mathcal{P}}(V)$ is defined as follows:
@@
Uncertainty sets of tree metrics.
Our key observation is that the computation of OT between input measures with tree metric does not depend on edges which have 00-length (i.e., $w_{e}=0$). We formally summarize it in Theorem 3.1.
@@
Let consider tree ${\mathcal{T}}_{1}$ as in Figure 1 and set $w_{e_{1}}=0$. Tree ${\mathcal{T}}_{2}$ is constructed from ${\mathcal{T}}_{1}$ as in Figure 1. Given two probability measures $\mu=a_{1}\delta_{r}+a_{2}\delta_{v_{1}}+a_{3}\delta_{v_{5}}$ and $\nu=b_{1}\delta_{v_{2}}+b_{2}\delta_{v_{4}}$ in ${\mathcal{T}}_{1}$ (i.e., $a_{1},a_{2},a_{3},b_{1},b_{2}\geq 0;a_{1}+a_{2}+a_{3}=1;b_{1}+b_{2}=1$), their corresponding measures in ${\mathcal{T}}_{2}$ are $\mu_{{\mathcal{T}}_{2}}=(a_{1}+a_{2})\delta_{r}+a_{3}\delta_{v_{5}}$ and $\nu_{{\mathcal{T}}_{2}}=\nu$ respectively. Then, we have
${\mathcal{W}}_{{\mathcal{T}}_{1}}(\mu,\nu)={\mathcal{W}}_{{\mathcal{T}}_{2}}(\mu_{{\mathcal{T}}_{2}},\nu_{{\mathcal{T}}_{2}}).$
@@
Let consider tree ${\mathcal{T}}_{1}$ as in Figure 1. Tree ${\mathcal{T}}_{3}$ is constructed from ${\mathcal{T}}_{1}$ as in Figure 1, and we set $w_{e_{a}}=w_{e_{b}}=0$. Given two probability measures $\mu=a_{1}\delta_{v_{1}}+a_{2}\delta_{v_{5}}$ and $\nu=b_{1}\delta_{v_{2}}+b_{2}\delta_{v_{4}}$ in ${\mathcal{T}}_{1}$ (i.e., $a_{1},a_{2},b_{1},b_{2}\geq 0;a_{1}+a_{2}=1;b_{1}+b_{2}=1$), their corresponding measures in ${\mathcal{T}}_{3}$ are $\mu_{{\mathcal{T}}_{3}}=a_{11}\delta_{v_{a}}+a_{12}\delta_{v_{1}}+a_{2}\delta_{v_{5}}$ where $a_{11},a_{12}\geq 0;a_{11}+a_{12}=a_{1}$ and $\nu_{{\mathcal{T}}_{3}}=\nu$ respectively. Then, regardless any split mass ($a_{11},a_{12}$) in $\mu_{{\mathcal{T}}_{3}}$ in ${\mathcal{T}}_{3}$ from original $\mu$ in ${\mathcal{T}}$, we have
${\mathcal{W}}_{{\mathcal{T}}_{1}}(\mu,\nu)={\mathcal{W}}_{{\mathcal{T}}_{3}}(\mu_{{\mathcal{T}}_{3}},\nu_{{\mathcal{T}}_{3}}).$
Therefore, we can add or delete edges with 00-length on the given tree ${\mathcal{T}}$ without changing the OT distance. More concretely,
  • (i) for edge deletion, we collapse edges with 00-length in ${\mathcal{T}}$ by merging the two corresponding vertices of those edges together (see Example 3.2).
  • (ii) for edge addition, we duplicate any vertex in ${\mathcal{T}}$ and connect them with 00-length edge (see Example 3.3).
These actions help to transform the given tree structure ${\mathcal{T}}$ into various tree structures, which play the fundamental role to construct uncertainty sets with diverse tree structures. Additionally, we further vary edge lengths of these tree structures to derive our novel uncertainty sets of tree metrics.
@@
(i) Constraints on individual edge. We consider an uncertainty for each edge length $\hat{w}_{e}$ of edge $e\in E$ in tree ${\mathcal{T}}(\hat{w})$. Specifically, we consider $\hat{w}_{e}$ belongs to some uncertainty interval $w_{e}-\alpha_{e}\leq\hat{w}_{e}\leq w_{e}+\beta_{e}$ around the edge weight $w_{e}$ in tree ${\mathcal{T}}$. In the vector form, this just means that
$w-\alpha\leq\hat{w}\leq w+\beta,$
with $\alpha,\,\beta\in\mathbb{R}^{|E|}_{+}$ satisfying $\alpha\leq w$ (i.e., edge weights are nonnegative). Thus, $w-\alpha$ and $w+\beta$ are respectively the lower and upper limits of the uncertainty interval for the vector of edge lengths.
@@
Closed-form expressions.
By building upon the proposed uncertainty sets of tree metrics and leveraging tree structure, we derive closed-form expressions for the max-min robust OT, similar to its counterpart standard OT (i.e., TW).
@@
Connection between two approaches.
We next draw a connection between two approaches for the robust OT for measures with noisy tree metric as follows:
@@
Assume that $\beta=\lambda\mathds{1}$. Then, we have
@@
Computational complexity.
From the closed-form expressions for $\textsf{RT}_{{\mathcal{U}}}$ in (7) and for $\textsf{RT}_{{\mathcal{U}}_{p}}$ in (10), the computational complexity of robust OT $\textsf{RT}_{{\mathcal{U}}}$ and $\textsf{RT}_{{\mathcal{U}}_{p}}$ for measures with noisy tree metric is linear to the number of edges in ${\mathcal{T}}$ (i.e., $\mathcal{O}(|E|)$), which is in the same order of computational complexity as the counterpart standard OT for measures with tree metric (i.e., TW) (Ba et al., 2011; Le et al., 2019). Recall that, in general, the max-min robust OT problem is hard and expensive to compute due to its non-convexity and non-smoothness (Paty & Cuturi, 2019; Lin et al., 2020), even for measures supported in one-dimensional space (Deshpande et al., 2019). [^5] Even with a given optimal ground metric cost, the computational complexity of max-min/min-max robust OT is in the same order as their counterpart standard OT (i.e., their objective function).
@@
Negative definiteness.
We next prove that the robust OT for measures with noisy tree metric is negative definite. Therefore, we can derive positive definite kernels built upon the robust OT.
@@
Metric property.
We end this section by showing that the robust OT is a metric.